OpenStudy (saifoo.khan):

Find the polynomial function with roots 11 and 2i. Steps please,

OpenStudy (anonymous):

oh my gosh seafood needs help? RARE....

OpenStudy (saifoo.khan):

lol^

OpenStudy (saifoo.khan):

YO?

OpenStudy (saifoo.khan):

chay-C?

OpenStudy (chaise):

I'm just trying to comprehend what you're asking, give me a sec.

OpenStudy (saifoo.khan):

Given..

OpenStudy (saifoo.khan):

(x-11)(x-2i)

OpenStudy (anonymous):

im hungry...

OpenStudy (saifoo.khan):

Batman, eat yourself.

OpenStudy (anonymous):

Heyy..since wen do bats eat them selves..i got a chocolate bar instead

OpenStudy (chaise):

What you're asking is: When you solve the polynomial, the answers roots are 11 and 2i. This means that the polynomials answers must be 11^2 and (2i)^2 Therefore, the polynomials answers must be 121 and -4 (x-121)(x+4)=0 x = +121 x = -4 This can't take the root of the negative, so you get 2i, you can take the root of 121. So you get 11. Expand that equation to get the polynomial. (x-121)(x+4) = x^2-117-484 Does this seem logical? I've never seen a question like this and this is more than likely wrong.

OpenStudy (saifoo.khan):

Alright Chocoman, try eating yourself.

OpenStudy (anonymous):

like i said..i cant eat my self..i cant wait till dinner..the whole house smells sooo good..i think ima eat my house

OpenStudy (chaise):

Oops - that should be x^2-117x-484. Sorry.

OpenStudy (chaise):

Does this seem logical and right?

OpenStudy (saifoo.khan):

umm... ummm.... nope. :(

OpenStudy (saifoo.khan):

why didn't u directly multiplied (x-11)(x-2i)

OpenStudy (anonymous):

where is the root -2i?

OpenStudy (saifoo.khan):

x = 11 ------------> x -11 = 0 x = 2i -------------> x - 2i = 0

OpenStudy (anonymous):

complex roots must occur in conjugate pairs; if 2i is a root, then -2i must also be a root; the poly for those two would be x^2+4=0

OpenStudy (anonymous):

eww...this is wat ima deal with in wat is this calculus? eww...

OpenStudy (chaise):

You've jumped a step - The question says. The polynomials ROOTS are 11 and 2i If you square both of them you get the direct answer x = 121 -----------------> x-121 x = -4 -----------------> x +4

OpenStudy (chaise):

This isn't calculus.

OpenStudy (saifoo.khan):

why u will square them?

OpenStudy (anonymous):

wat is this..i wanna prepare...

OpenStudy (anonymous):

this is basic algebra

OpenStudy (anonymous):

x^2+4=(x-2i)(x+2i)

OpenStudy (saifoo.khan):

yes, right @mandolio.. then/

OpenStudy (saifoo.khan):

guys, i get till where mandolio ended.. then wht should i dp?

OpenStudy (saifoo.khan):

do*

OpenStudy (anonymous):

If the poly is third degree then take your pick: (x+11)(x-2i)(x+2i)=0 (x+11)(x^2+4)=0 x^3+11x^2+4x+44=0

OpenStudy (vishweshshrimali5):

(x-11)(x-2i) this is the answer

OpenStudy (saifoo.khan):

Alright, thanks!! :)

OpenStudy (vishweshshrimali5):

@ Mandolino I think the roots of ur eq. are -11, 2i, -2i

OpenStudy (vishweshshrimali5):

but the roots re. are 11, 2i

OpenStudy (anonymous):

check this out-- 2i by itself can't be a root!!!!! http://en.wikipedia.org/wiki/Complex_conjugate_root_theorem

OpenStudy (anonymous):

First, chaise, you're way off base with squaring the roots first. Second, a polynomial with REAL coefficients must have complex conjugate solutions (if at least one complex solution exists). The problem doesn't say the polynomial must have real coefficients. (x-11)(x-2i) = x^2 - 2ix - 11x + 22i = x^2 - (11+2i)x +22i That is a polynomial. Its roots are 11 and 2i. QED